OpenStudy (anonymous):

the length of a rectangle is 5 cm and the width is 4 cm. Find the length of the diagonal

OpenStudy (bahrom7893):

25+16=d^2

OpenStudy (anonymous):

4^2 + 5^2 = c^2

OpenStudy (bahrom7893):

d^2=41 d=sqrt(41)

OpenStudy (anonymous):

√41